# How Do Recommendation Models Amplify Popularity Bias? An Analysis from the Spectral Perspective

### v3

@@ Appendix B Additional Experiments
B.1 Long-tailed Distribution and Bias Amplification in Recommendations
@@
Figure 7: Long-tailed distribution of item popularity in recommendation datasets.
Figure 8: Illustration of popular bias amplification: We divide items into five groups according to their popularity as recent work [68], and focus on the most popular group. The chart displays four bars, representing the ratio of the items in the most popular group, the percentage of interactions originating from these items in the training set, and the percentages of these items appearing in recommendations from MF and LightGCN, respectively.
Table 4: The proportion of interactions of the top 20% of popular items in the total number of interactions.
@@
Table 5: The value of $\sigma_{1}^{2}$ and $r_{max}$ in different datasets and recommendation models.
@@
Table 6: Comparison between the actual spectral norm and the estimated approximation.
Datasets MSE BCE
$||{\bf{UV^{\top}}}||_{2}^{2}$ $||{\bf{U(V^{\top}{{\bf{\tilde{q}}}_{1}})}}||^{2}$ $||{\bf{UV^{\top}}}||_{2}^{2}$ $||{\bf{U(V^{\top}{{\bf{\tilde{q}}}_{1}})}}||^{2}$
Figure 9: Performance comparison across different embedding dimensions in the Movielens and Douban datasets.
Table 7: Running time comparison (s / Epoch).
Movielens Douban
MF 0.177 2.098
ReSN 0.181 2.124
ReSN-Direct 649 59239
Speedup Ratio 3585 27890
B.2 Comparison between $\sigma_{1}^{2}$ and $r_{max}$
Table 5 presents the values of $\sigma_{1}^{2}$ and $r_{max}$ on three benchmark datasets and different backbone models. It can be observed that in multiple actual datasets, $\sigma_{1}^{2}$ is significantly larger than $r_{max}$, exceeding $100$ times and more. Combining the bounds given by Eq.(1), ($\cos({\bf{r}},{{\bf{q}}_{1}})\geq\frac{{\sigma_{1}^{2}}}{{{r_{\max}}\sqrt{\zeta(2\alpha)}}}\sqrt{1-\frac{{{r_{\max}}(\zeta(\alpha)-1)}}{{\sigma_{1}^{2}}}}$), even if the data is not markedly skewed, i.e., $\alpha$ is not very large and $\zeta(\alpha)$ is not very close to 1, there is still a significant similarity between item popularity vector $\bf{r}$ and the principal singular vector $\bf{q}_{1}$ due to the considerable ratio between $\sigma_{1}^{2}$ and $r_{max}$. This observation also helps to explain the prevalence of popularity bias memorization effect in recommendation models and datasets.
B.3 Validation on Approximation
We computed the ideal value of $||{\mathbf{U}\mathbf{V}^{\top}}||_{2}^{2}$, as well as the estimated $\frac{||{\bf{U}}{{\bf{V}}^{\top}}{\bf{V}}{{\bf{U}}^{\top}}{\bf{e}}||^{2}}{{||{\bf{U}\bf{V}^{\top}{\bf{e}}}||^{2}}}$ from ReSN. Here we well-training the MF model with two losses on three datasets. The results are shown in the Table 6. According to the table, we found that the actual spectral norms and our approximate estimates are very close across diverse losses and datasets.This indicates that the singular vector $\bf{\tilde{q}}_{1}$ obtained through $\frac{\bf{U}\bf{V}^{\top}\bf{e}}{||\bf{U}\bf{V}^{\top}\bf{e}||}$, serves as an accurate surrogate for the true value of $\bf{q}_{1}$. Therefore, the estimated regularization term is a accurate surrogate for the spectral norm $||{\bf{U}\bf{V}^{\top}}||^{2}_{2}$ which validates the accuracy and rationality of the proposed method.
B.4 Performance with Diverse Embedding sizes
@@
B.5 Efficiency Study
To further validate the effectiveness of our acceleration strategy, we test the running time per epoch of ReSN and the original brute-force strategy employed to compute the gradient of the spectral norm. Also, we present the baseline MF for comparison. The results are presented in Table 7. As can be seen, our acceleration strategy achieves over 3500 and 27000 times impressive speed-up in both datasets, respectively. Moreover, compared with MF, our ReSN does not incur much computational overhead.
Table 8: Dataset statistics.
@@ B.5 Efficiency Study
Table 9: Notations in this paper.
@@
$\bf{Y}$ the observed interaction matrix
@@
$\bf{u}_{u}$,$\bf{v}_{i}$ embedding vector of user $u$ and item $i$
$\bf{U}$,$\bf{V}$ embedding matrices for all users and items
$\bf{\hat{Y}}$ predicted matrix of all user-item pairs,i.e., $\bf{\hat{Y}}=\mu(\bf{U}\bf{V}^{\top})$
$\sigma_{k}$,$\bf{p}_{k}$,$\bf{q}_{k}$ the $k$-th largest singular value and its corresponding left and right singular vector of $\bf{\hat{Y}}$
@@
$\bf{e}$ a $n$-dimension vector filled with ones
@@ C.1 Datasets
  • Movielens-1M [63]: Movielens is the widely used dataset from [63] and is collected from MovieLens [^2] https://movielens.org/. We use the version of 1M. We transform explicit data into implicit feedback by treating all user-item ratings as positive interactions.
  • Douban [50]: This dataset is collected from a popular review website Douban [^3] https://www.douban.com/ in China. We transform explicit data into implicit using the same method as applied in Movielens.
  • Globo [20]: This dataset is a popular dataset collected from the news recommendation website Globo.com [^4] http://g1.globo.com/.
  • Yelp2018 [27] & Gowalla [26]: Gowalla is the check-in dataset obtained from Gowalla and Yelp2018 is from the 2018 edition of the Yelp challenge, containing Yelp’s business reviews and user data. For a fair comparison, these two datasets are used exactly the same as [27] used.
  • Yahoo!R3 [38] & Coat [49]: These two datasets are obtained from the Yahoo music and Coat shopping recommendation service, respectively. Both datasets contain a training set of biased rating data collected from normal user interactions and a test set of unbiased rating data containing user ratings on randomly selected items. The rating data are translated to implicit feedback,i.e., interactions with ratings larger than 3 are regarded as positive samples.
@@
C.2 Implementation Details
We implement ReSN in Tensorflow [1] and the initialization is unified with Xavier [24]. We optimize all models with Adam [31]. A grid search is conducted to confirm the optimal parameter setting for each model. To be more specific, learning rate is searched in $\{1e^{-2},1e^{-3},2e^{-4}\}$, weight decay in $\{1e^{-7},1e^{-6},1e^{-5},1e^{-4},1e^{-3}\}$. As for the backbone of LightGCN, we utilize three layers of graph convolution network to obtain the best results, with or without using dropout to prevent over-fitting. For ReSN, the coefficient of regularizer $\beta$ is tuned in the range of $\{1e^{-4},1e^{-3},1e^{-2},1e^{-1},5e^{-1},1.0,5.0\}$. For compared methods, we closely refer to configurations provided in their respective publications to ensure their optimal performance.

### v6

@@ Appendix B Additional Experiments
B.1. Long-tailed Distribution and Bias Amplification in Recommendations
@@
Figure 5. Long-tailed distribution of item popularity in recommendation datasets.
Figure 6. Illustration of popular bias amplification: We divide items into five groups according to their popularity as recent work (Zhang et al., 2021), and focus on the most popular group. The chart displays four bars, representing the ratio of the items in the most popular group, the percentage of interactions originating from these items in the training set, and the percentages of these items appearing in recommendations from MF and LightGCN, respectively.
Table 8. The proportion of interactions of the top 20% of popular items in the total number of interactions.
@@
Table 9. The value of $\sigma_{1}^{2}$ and $r_{max}$ in different datasets and recommendation models.
@@
B.2. Comparison between $\sigma_{1}^{2}$ and $r_{max}$
Table 9 presents the values of $\sigma_{1}^{2}$ and $r_{max}$ on three benchmark datasets and different backbone models. It can be observed that in multiple actual datasets, $\sigma_{1}^{2}$ is significantly larger than $r_{max}$, exceeding $100$ times and more. Combining the bounds given by Eq.(1), ($\cos({\mathbf{r}},{{\mathbf{q}}_{1}})\geq\frac{{\sigma_{1}^{2}}}{{{r_{\max}}\sqrt{\zeta(2\alpha)}}}\sqrt{1-\frac{{{r_{\max}}(\zeta(\alpha)-1)}}{{\sigma_{1}^{2}}}}$), even if the data is not markedly skewed, i.e., $\alpha$ is not very large and $\zeta(\alpha)$ is not very close to 1, there is still a significant similarity between item popularity vector $\mathbf{r}$ and the principal singular vector $\mathbf{q}_{1}$ due to the considerable ratio between $\sigma_{1}^{2}$ and $r_{max}$. This observation also helps to explain the prevalence of popularity bias memorization effect in recommendation models and datasets.
Figure 7. Performance comparison across different embedding dimensions in the Movielens and Douban datasets.
B.3. Performance with Diverse Embedding sizes
@@
Table 10. Dataset statistics.
@@
Table 11. Notations in this paper.
@@
$\mathbf{Y}$ the observed interaction matrix
@@
$\mathbf{u}_{u}$,$\mathbf{v}_{i}$ embedding vector of user $u$ and item $i$
$\mathbf{U}$,$\mathbf{V}$ embedding matrices for all users and items
$\mathbf{\hat{Y}}$ predicted matrix of all user-item pairs,i.e., $\mathbf{\hat{Y}}=\mu(\mathbf{U}\mathbf{V}^{\top})$
$\sigma_{k}$,$\mathbf{p}_{k}$,$\mathbf{q}_{k}$ the $k$-th largest singular value and its corresponding left and right singular vector of $\mathbf{\hat{Y}}$
@@
$\mathbf{e}$ a $n$-dimension vector filled with ones
@@ C.1. Datasets
  • Movielens-1M (Yu and Qin, 2020): Movielens is the widely used dataset from (Yu and Qin, 2020) and is collected from MovieLens [^2] https://movielens.org/. We use the version of 1M. We transform explicit data into implicit feedback by treating all user-item ratings as positive interactions.
  • Douban (Song et al., 2019): This dataset is collected from a popular review website Douban [^3] https://www.douban.com/ in China. We transform explicit data into implicit using the same method as applied in Movielens.
  • Globo (de Souza Pereira Moreira et al., 2018): This dataset is a popular dataset collected from the news recommendation website Globo.com [^4] http://g1.globo.com/.
  • Yelp2018 (He et al., 2020) & Gowalla (He et al., 2017): Gowalla is the check-in dataset obtained from Gowalla and Yelp2018 is from the 2018 edition of the Yelp challenge, containing Yelp’s business reviews and user data. For a fair comparison, these two datasets are used exactly the same as (He et al., 2020) used.
  • Yahoo!R3 (Marlin and Zemel, 2009) & Coat (Schnabel et al., 2016): These two datasets are obtained from the Yahoo music and Coat shopping recommendation service, respectively. Both datasets contain a training set of biased rating data collected from normal user interactions and a test set of unbiased rating data containing user ratings on randomly selected items. The rating data are translated to implicit feedback,i.e., interactions with ratings larger than 3 are regarded as positive samples.
@@
C.2. Implementation Details
We implement ReSN in Tensorflow (Abadi et al., 2016) and the initialization is unified with Xavier (Glorot and Bengio, 2010). We optimize all models with Adam (Kingma, 2014). A grid search is conducted to confirm the optimal parameter setting for each model. To be more specific, learning rate is searched in $\{1e^{-2},1e^{-3},2e^{-4}\}$, weight decay in $\{1e^{-7},1e^{-6},1e^{-5},1e^{-4},1e^{-3}\}$. As for the backbone of LightGCN, we utilize three layers of graph convolution network to obtain the best results, with or without using dropout to prevent over-fitting. For ReSN, the coefficient of regularizer $\beta$ is tuned in the range of $\{1e^{-4},1e^{-3},1e^{-2},1e^{-1},5e^{-1},1.0,5.0\}$. For compared methods, we closely refer to configurations provided in their respective publications to ensure their optimal performance.